If f(x) = 7x3 − 8x + 2, then f(8) = ___.
Answer by checkley77(12569)

If f(x) = 7x3 − 8x + 2, then f(8)
f(8)=7*8^3-8*8+2
f(8)=7*512-64+2
f(8)=3,584-62
f(8)=3,522 ans.